How many grams of Cu are needed to react with 105.2g of Ag
1
Expert's answer
2020-05-27T13:34:11-0400

"2Ag^+ + Cu = Cu(2+) + 2 Ag"

"n(Ag) = 105.2\/107.87 mol = 0.975 mol"

"n(Cu) = n(Ag)\/2 = 0.4875 mol => m(Cu) = 0.4875*63.55 = 30.98 g"